OpenStudy (anonymous):

If y varies inversely as x and y is 2.5 when x is 6, what is y when x is 4?

OpenStudy (anonymous):

y is proportional to (1/x) Let y=c/x where is c is a constant Given: x=6 --> y=2.5 Plug these values into the equation 2.5 = c/6 c= 6 * 2.5 = 15 So, y= 15/x When x=4, y=15/4 = 3.75

OpenStudy (anonymous):

Try this question in Math.